OpenStudy (anonymous):

What is the next number in the pattern: 1, -2, 5, 2, 9, 6, 13, …?

OpenStudy (anonymous):

answers: -6 20 22 10

ganeshie8 (ganeshie8):

see the pattern : -3, +7...

ganeshie8 (ganeshie8):

1-3 = -2 -2+7 = 5 5-3 = 2 2+7 = 9 9-3 = 6 6 +7 = 13 13-? = ??

OpenStudy (anonymous):

it would be 10

ganeshie8 (ganeshie8):

Yes !

ganeshie8 (ganeshie8):

good work :)

OpenStudy (anonymous):

Thank you could you help me another one?

ganeshie8 (ganeshie8):

yea sure

OpenStudy (anonymous):

What is the 8th number in the pattern: 20, 12, 4, -4, …?

ganeshie8 (ganeshie8):

this is just a arithmetic sequence, a = 20, d = ?

ganeshie8 (ganeshie8):

\(t_n = a +(n-1)d\) \(t_8 = 20 +(8-1)d\)

ganeshie8 (ganeshie8):

if u knw d, u can plug in above, and find the 8th term \(t_8\)

OpenStudy (anonymous):

d is 12 right?

ganeshie8 (ganeshie8):

not exactly.. d = next term - present term = 12 - 20 = ?

OpenStudy (anonymous):

-8

ganeshie8 (ganeshie8):

yes plugin above and find \(t_8\)

OpenStudy (anonymous):

ok I got -3.375 but that isn't correct

ganeshie8 (ganeshie8):

\(t_8 = 20 + (8-1)(-8)\) \( = 20 + (7)(-8)\) \( = 20 - 56\) \( = ?\)

OpenStudy (anonymous):

-36! thank you again

ganeshie8 (ganeshie8):

np :)